From: STOSC::SMM "On-Duty SMM x5074" 28-JAN-1993 23:35:50.82 To: @[SMM.DIS]HST_STAT CC: Subj: HST STATUS REPORT, 01/27 - 01/28 HUBBLE SPACE TELESCOPE DAILY REPORT #811 PERIOD COVERED: 0800 EST 01/27/93 - 0800 EST 01/28/93 Daily Status Report as of 028/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completion of the Zero-Gyro (On-Orbit) Test 026/1200Z through 028/0000Z. (See Operations Note 2.4). 1.2 Completed S/C YSAFE (Safe Mode Recovery of HST Sci. Inst./FOS) The Safe Mode recovery operations were successful. The FOS appears to be normal. 1.3 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.4 Completed GHRS/2 3816 (White Dwarf Stars) The High Resolution Spectrograph was used to make ultraviolet observations of the white dwarf star G358. The guide star acquisition was made in course track. The Return-to-Brightness spiral target acquisition located the target on the first map. The peakup was executed without problems. The conformation map showed that the target was well centered. 1.5 Completed FOS/RD 3296 (Super Nova Remnants and Nucleo Synthesis) The Faint Object Spectrograph was used on both the Red and Blue sides to make spectral observations of the supernova remnant NGC-4449-SN. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 4 Successful : 4 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Significant HSTARs: None 2.4 Operations Note: The Zero Gyro On-Orbit Test was successfully completed on Day 92.026-027. Commands to activate the Zero Gyro software were executed at approximately 026/1300Z, while the vehicle was in Inertial Hold during Orbit Day. The vehicle remained under Zero Gyro control until 027/0700Z, when Software Sunpoint safemode commands were executed as planned, to begin the recovery back to Normal Mode. During this recovery, the Aperture Door was successfully commanded open. All performance during Zero Gyro control, including the Free Drift during Night and Sun Capture during Orbit Day, was nominal. All test objectives were met. Normal science operations was resumed at approximately 028/0000Z. A NSSC-I Health and Safety Load was built and not loaded, because it was not needed. The Faint Object Spectrograph, Faint Object Camera, and the High Resolution Spectrograph were all successfully recovered from Safe Mode. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 3296 (Supernova Remnants and Nucleo Synthesis) 3.2 Complete FOS/RD 4206 (Spectroscopy of Super Metal Poor Galaxies) 3.3 Complete WFC/PC 4207 (Photometric Calibration Monitor) 3.4 Complete FOS/BL 4698 (FOS Dark Monitoring) 3.5 Complete PC 4140 (WF/PC Calibration:Rapid Internal Monitor - Part 2) 3.6 Uplink Updated Ephemeris Tables, Per MOSES 2871 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-